Factorize x⁴ - 16
1
Expert's answer
2021-03-23T08:12:50-0400
"x\u2074 - 16=(x^2)^2-(2^2)^2."

We know that


"a^2-b^2=(a-b)(a+b)."

Thus, putting


"a=x^2, b=2^2,"

we have


"x\u2074 - 16=(x^2)^2-(2^2)^2=(x^2-2^2)(x^2+2^2)."

Apply the rule for "x^2-2^2":


"(x^2-2^2)=((x-2)(x+2))."

Therefore,


"x\u2074 - 16=(x-2)(x+2)(x^2+4)"